# Diverse dark matter profiles in fire dwarfs: black holes, cosmic rays and the cusp-core enigma

### v1

@@ Abstract
pubyear: 2024pagerange: Diverse dark matter profiles in fire dwarfs: black holes, cosmic rays and the cusp-core enigmaDiverse dark matter profiles in fire dwarfs: black holes, cosmic rays and the cusp-core enigma
@@ 1 Introduction
However, there has also been a large body of theoretical work focusing on improving the baryonic physics modelling which may largely alleviate the ‘dwarf galaxy problems’ observed in pure CDM simulations. In particular, star formation suppression from reionization could significantly decrease the number of bright dwarf satellites (e.g. Efstathiou, 1992; Okamoto et al., 2008; Fitts et al., 2017; Katz et al., 2020) thereby resolving the missing satellite problem. Furthermore, stellar feedback (e.g. Navarro et al., 1996; Governato et al., 2010; Parry et al., 2012; Brooks & Zolotov, 2014; Hopkins et al., 2014, 2018; Chan et al., 2015; Kimm et al., 2015; Emerick et al., 2018; Smith et al., 2018, 2019; Gutcke et al., 2021; Jackson et al., 2024) may significantly decrease the central dark matter densities of dwarf galaxies. In particular, it has been demonstrated that cyclic supernova (SN) bursts play an important role in creating cores in dwarf galaxies (e.g. Pontzen & Governato, 2012) and that late-time star formation is crucial for maintaining these cores (e.g. Read et al., 2019; Muni et al., 2024).
@@
These tantalising observations have motivated theorists to investigate this largely unexplored regime. Analytical models point towards favourable AGN energetics in the dwarf regime (compared to stellar feedback, see e.g. Dashyan et al., 2018) and suggest that AGN activity may be able resolve all of the remaining ‘dwarf galaxy problems’ (Silk, 2017). However, hydrodynamical simulations have painted a more complex picture. AGN activity in dwarf galaxies may be significantly suppressed by stellar feedback evacuating gas from the central region (e.g. Dubois et al., 2015; Anglés-Alcázar et al., 2017b; Habouzit et al., 2017; Trebitsch et al., 2018; Hopkins et al., 2022a; Byrne et al., 2023a) and black holes (BHs) may be wandering in dwarf galaxies due to their shallow potential wells (e.g. Bellovary et al., 2021; Ma et al., 2021; Sharma et al., 2022; Beckmann et al., 2023) further decreasing their accretion efficiency. The modelling of BH growth is another crucial aspect as the fiducial Bondi model suppresses the growth of low-mass BHs compared to gas-supply-limited or torque-driven schemes due to its strong dependence on BH mass (e.g. Anglés-Alcázar et al., 2013, 2015; Koudmani et al., 2022; Wellons et al., 2023; Gordon et al., 2024). The cosmic environment also has a crucial influence on AGN activity in the dwarf regime with minor mergers triggering AGN episodes whilst long-term residence in dense environments is detrimental to BH growth (e.g. Kristensen et al., 2021).
@@
Despite these caveats, various groups have identified physical regimes where AGN in dwarfs can accrete efficiently in accordance with observed samples. These dwarf AGN significantly influence their host galaxies by driving powerful outflows (e.g. Koudmani et al., 2019, 2022) and suppressing star formation (e.g. Barai & de Gouveia Dal Pino, 2019; Sharma et al., 2020, 2022, 2023; Koudmani et al., 2021, 2022; Wellons et al., 2023; Arjona-Galvez et al., 2024). Given this potentially significant impact by AGN on the baryon cycle in dwarf galaxies, could AGN feedback also impact the dark matter distributions in dwarfs as predicted by analytical models?
@@
The interaction between AGN feedback, star formation, and dark matter in dwarf galaxies has not been thoroughly explored, though some trends have been noted, such as AGN mildly suppressing central dark matter densities (see Koudmani et al., 2022; Arjona-Galvez et al., 2024), hinting at a possible role of AGN in driving cusp-to-core transformations in dwarfs. However, the central density suppressions by the AGN in these simulations were mostly minor and did not change the qualitative nature of the central profiles. This is likely due to the effective equation of state (Springel & Hernquist, 2003) employed for modelling the interstellar medium (ISM) in these simulations, which leads to smoother star formation histories alongside less bursty AGN feedback at low redshifts, often operating as ‘maintenance-mode’ feedback (also see discussion in Koudmani et al., 2022). Recently, Arora et al. (2024) examined the role of BH feedback in core formation within the NIHAO simulations and found negligible effects in low-mass systems, largely due to inefficient BH growth in this mass regime, constrained by the Bondi accretion scheme. In contrast, they found significant impacts of BH feedback on core formation in more massive galaxies, aligning with earlier studies that show efficient AGN activity can create cores in these galaxies (e.g. Martizzi et al., 2013; Peirani et al., 2017; Macciò et al., 2020).
@@
It is therefore timely to investigate whether efficient AGN activity can induce cusp-to-core transformations in dwarf galaxies and potentially account for the observed diversity in dwarf rotation curves. Clearly, within this context it is crucial to model the multi-phase nature of the ISM but also to consider the role of non-thermal components such as magnetic fields or cosmic rays (CRs) (e.g. Uhlig et al., 2012; Booth et al., 2013; Hanasz et al., 2013; Salem & Bryan, 2014; Pakmor et al., 2016; Farber et al., 2018; Holguin et al., 2019; Dashyan & Dubois, 2020; Martin-Alvarez et al., 2020; Steinwandel et al., 2020). In particular, it has been shown that CRs may drive core formation (e.g. Hopkins et al., 2020b; Martin-Alvarez et al., 2023) though their impact in the dwarf regime remains controversial as gas-rich mergers may re-establish cores in these systems (see discussion in Martin-Alvarez et al., 2023). What is more, it has been found that CRs may significantly boost the efficiency of AGN feedback (e.g. Su et al., 2021, 2024; Byrne et al., 2023b; Wellons et al., 2023) by suppressing cold gas accretion from the CGM due to enhanced CR pressure support. Therefore, it is important to also assess the role of CRs when investigating cusp-core transformations with AGN.
@@ 2.1 General set-up
Our simulation suite is based on a set of $3$ initial conditions (ICs) of dwarf haloes selected from the fire ICs set, spanning a range of halo masses in the ‘classical dwarf’ regime. We select two low-mass dwarfs: m10q and m10y, with $z=0$ halo masses of $M_{\mathrm{vir}}=8\times 10^{9}$ $\mathrm{M_{\odot}}$ and $M_{\mathrm{vir}}=1.4\times 10^{10}$ $\mathrm{M_{\odot}}$, respectively. Our third halo, m10z, is an intermediate-mass dwarf with $z=0$ halo mass $M_{\mathrm{vir}}=3.5\times 10^{10}$ $\mathrm{M_{\odot}}$. All of these ICs were generated using the music IC generator (Hahn & Abel, 2011) and have been thoroughly explored with the fire-2 model (Hopkins et al., 2018). In the standard fire-2 runs, m10q has a quiescent growth history and is relatively isolated, m10y forms early and has a large core, and m10z is an ultra-diffuse galaxy.
To simulate these systems, we use the gizmo hydrodynamics+gravity code, employing the mesh-free finite mass method (Lanson & Vila, 2008a, b; Gaburov & Nitadori, 2011; Hopkins, 2015, 2017) that tracks constant-mass gas resolution elements in a Lagrangian fashion. Our simulations include both baryons and dark matter, with mass resolutions of $M_{\mathrm{gas}}=262$ $\mathrm{M_{\odot}}$ and $M_{\mathrm{DM}}=1303$ $\mathrm{M_{\odot}}$, respectively. The gravitational softening for the dark matter, stars and BH particles is set to $\epsilon_{\mathrm{soft}}=0.0625\ \mathrm{ckpc}\,h^{-1}$, whereas the softening for the gas resolution elements is set adaptively. This allows us to easily resolve core formation as demonstrated in previous work (e.g. Chan et al., 2015; Hopkins et al., 2018).
@@ 2.2 Galaxy formation model
For the galaxy formation physics, we use the newly updated fire-3 model (Hopkins et al., 2023a) which tracks the state of the multi-phase ISM and multiple forms of stellar feedback, including feedback from SNee [^1] We couple the SN momentum following Hopkins et al. (2023a) and do not yet include the updates to the momentum coupling from Hopkins (2024). Though we note that these changes mainly affect high-mass galaxies at low resolution, whilst the impact on high-resolution dwarfs is very limited. of Type Ia and II, stellar winds from OB and AGB stars as well as multi-wavelength photo-heating and radiation pressure. Star formation occurs in gas that is molecular, dense, self-gravitating, self-shielding, and Jeans-unstable. All simulations include magnetic fields, using the magneto-hydrodynamics methods from Hopkins & Raives (2016) and Hopkins (2016). The fire-3 model also has new optional physics modules compared to the previous fire (Hopkins et al., 2014) and fire-2 (Hopkins et al., 2018) models. Two of the most important optional features are a novel sub-grid BH model, including seeding, accretion and feedback, as well as the inclusion of CR physics. To separate the impacts of CRs and BHs, we perform four simulations for each set of ICs, including runs without CRs or BHs, with both CRs and BHs and with only CRs or BHs, respectively. For the runs without BHs but with CRs, no BHs are seeded and the CRs only stem from SNe. For the runs without CRs and with BHs, we incorporate all AGN feedback channels except for CR injection and SNe do not inject CRs either. See Table 1 for an overview of the simulation suite.
@@
Table 1: Overview of dwarf zoom-in simulation suite. We list the simulation names (first column), ICs (second column), and whether BHs (third column) and CRs (fourth column) are included. We also provide the central dark matter densities at $r=150$ pc (fifth column), the ratios between the central dark matter density for the hydro runs and dark matter only simulations (sixth column), and the shape of the central dark matter density profile (seventh column) as shown in Figure 8. All density estimators are rounded to two significant figures.
@@
Simulation name ICs BHs? CRs? $\boldsymbol{\rho_{\mathrm{DM,inner}}\ [10^{8}\ \mathrm{M_{\odot}}\,\mathrm{kpc^{-3}}]}$ $\boldsymbol{\rho_{\mathrm{DM,inner}}/\rho_{\mathrm{DMO,inner}}}$ Central dark matter profile
m10q m10q No No 0.73 0.47 weak core
m10qAGN m10q Yes No 0.42 0.27 strong core
m10qCR m10q No Yes 1.4 0.92 weak cusp
m10qCR+AGN m10q Yes Yes 1.5 0.99 weak cusp
m10y m10y No No 0.20 0.062 strong core
m10yAGN m10y Yes No 0.19 0.058 strong core
m10yCR m10y No Yes 0.83 0.26 weak core
m10yCR+AGN m10y Yes Yes 0.74 0.23 weak core
m10z m10z No No 0.22 0.10 strong core
m10zAGN m10z Yes No 0.069 0.032 strong core
m10zCR m10z No Yes 0.11 0.053 strong core
m10zCR+AGN m10z Yes Yes 0.66 0.31 weak core
@@
The modelling of BHs and CRs remains highly uncertain. Several studies have investigated the impact of different modelling assumptions for CRs (e.g. Hopkins et al., 2020b, 2023b; Butsky et al., 2023) and BHs (Anglés-Alcázar et al., 2017b, 2021; Su et al., 2021, 2024; Byrne et al., 2023a, b; Wellons et al., 2023) within the fire model. For our study, we base our BH model on the extensive study by Wellons et al. (2023) who identified a modelling space that allows for efficient BH feedback whilst still matching observational constraints, in particular with regards to star formation distributions. We summarise these BH modelling choices in Section 2.2.1 and point the interested reader to Wellons et al. (2023) for more details and an in-depth parameter study of alternatives to our choices. For the CR modelling, we employ the numerically-efficient subgrid CR model introduced in Hopkins et al. (2023b) allowing us to explore several different IC set-ups to $z=0$ whilst minimising the computational overhead costs that would arise from explicit CR transport. We summarise the key characteristics of the CR subgrid model in Section 2.2.2 and refer the interest reader to Hopkins et al. (2023b) for a more detailed overview of this approach.
@@ 2.2.1 BH modelling
Once a BH particle is present, it grows via accretion by draining gas from the nearest $N\sim 256$ gas cell neighbours. Wellons et al. (2023) parameterized the accretion rate as $\dot{M}_{\mathrm{acc}}\equiv\eta_{\mathrm{acc}}M_{\mathrm{gas}}\Omega$, where $\eta_{\mathrm{acc}}$ is the accretion efficiency, $M_{\mathrm{gas}}$ is the gas mass within the BH kernel, and $\Omega=\sqrt{GM_{\mathrm{tot}}/R}$ is the orbital frequency determined at the kernel size $R$, based on the total mass enclosed with $R$, $M_{\mathrm{tot}}$. fire-3 modulates the BH accretion rate via a gas reservoir that represents an accretion disc from which the BH particle grows at a mass growth rate $\dot{M}_{\mathrm{BH}}=M_{\mathrm{\alpha disc}}/t_{\mathrm{dep}}$, where $M_{\mathrm{\alpha disc}}$ is the mass of the accretion disc and $t_{\mathrm{dep}}=42~{}\mathrm{Myr}\ (1+M_{\mathrm{BH}}/M_{\mathrm{\alpha disc}})^{0.4}$ is the depletion time.
@@
We set $\epsilon_{\mathrm{jet}}=0.35$ which leads to an effective base opening angle of $\sim 7\degree$. The angle weighting is applied with respect to the total gas angular momentum accreted by the BH, $J_{\mathrm{gas}}$, which is measured in the centre-of-mass frame of the BH – accretion disc sink particle. This allows for AGN energy injection without significantly disrupting the small-scale gas supply in the central region which favours repeated bursts and/or continuous AGN activity allowing us to probe the ‘physically interesting’ regime for AGN influencing cusp to core transformations. As discussed in Su et al. (2021), the large-scale collimation properties of the jet are not hugely sensitive to the small-scale opening angle since the feedback energy chooses the path of least resistance effectively recollimating even for large opening angles, also see Koudmani et al. (2019). Long-range radiation transport and CRs are treated via the LEBRON approximation and injected isotropically (see Hopkins et al., 2020a, 2023b).

### v2

# Diverse dark matter profiles in fire dwarfs: black holes, cosmic rays and the cusp–core enigma

@@ Abstract
pubyear: 2024pagerange: Diverse dark matter profiles in fire dwarfs: black holes, cosmic rays and the cusp–core enigmaDiverse dark matter profiles in fire dwarfs: black holes, cosmic rays and the cusp–core enigma
@@ 1 Introduction
However, there has also been a large body of theoretical work focusing on improving the baryonic physics modelling which may largely alleviate the ‘dwarf galaxy problems’ observed in pure CDM simulations. In particular, star formation suppression from reionization could significantly decrease the number of bright dwarf satellites (e.g. Efstathiou, 1992; Okamoto et al., 2008; Fitts et al., 2017; Katz et al., 2020) thereby resolving the missing satellite problem. Furthermore, stellar feedback (e.g. Navarro et al., 1996; Governato et al., 2010; Parry et al., 2012; Brooks & Zolotov, 2014; Hopkins et al., 2014, 2018; Chan et al., 2015; Kimm et al., 2015; Emerick et al., 2018; Smith et al., 2018, 2019; Gutcke et al., 2021; Jackson et al., 2024) may significantly decrease the central dark matter densities of dwarf galaxies. In particular, it has been demonstrated that cyclic supernova (SN) bursts play an important role in creating cores in dwarf galaxies (e.g. Pontzen & Governato, 2012) and that late-time star formation is crucial for maintaining these cores (e.g. Read et al., 2019; Muni et al., 2025).
@@
These tantalising observations have motivated theorists to investigate this largely unexplored regime. Analytical models point towards favourable AGN energetics in the dwarf regime (compared to stellar feedback, see e.g. Dashyan et al., 2018) and suggest that AGN activity may be able resolve all of the remaining ‘dwarf galaxy problems’ (Silk, 2017). However, hydrodynamical simulations have painted a more complex picture. AGN activity in dwarf galaxies may be significantly suppressed by stellar feedback evacuating gas from the central region (e.g. Dubois et al., 2015; Anglés-Alcázar et al., 2017b; Habouzit et al., 2017; Trebitsch et al., 2018; Hopkins et al., 2022a; Byrne et al., 2023) and black holes (BHs) may be wandering in dwarf galaxies due to their shallow potential wells (e.g. Bellovary et al., 2021; Ma et al., 2021; Sharma et al., 2022; Beckmann et al., 2023) further decreasing their accretion efficiency. The modelling of BH growth is another crucial aspect as the fiducial Bondi model suppresses the growth of low-mass BHs compared to gas-supply-limited or torque-driven schemes due to its strong dependence on BH mass (e.g. Anglés-Alcázar et al., 2013, 2015; Koudmani et al., 2022; Wellons et al., 2023; Gordon et al., 2024). The cosmic environment also has a crucial influence on AGN activity in the dwarf regime with minor mergers triggering AGN episodes whilst long-term residence in dense environments is detrimental to BH growth (e.g. Kristensen et al., 2021).
@@
Despite these caveats, various groups have identified physical regimes where AGN in dwarfs can accrete efficiently in accordance with observed samples. These dwarf AGN significantly influence their host galaxies by driving powerful outflows (e.g. Koudmani et al., 2019, 2022) and suppressing star formation (e.g. Barai & de Gouveia Dal Pino, 2019; Sharma et al., 2020, 2022, 2023; Koudmani et al., 2021, 2022; Wellons et al., 2023; Arjona-Gálvez et al., 2024). Given this potentially significant impact by AGN on the baryon cycle in dwarf galaxies, could AGN feedback also impact the dark matter distributions in dwarfs as predicted by analytical models?
@@
The interaction between AGN feedback, star formation, and dark matter in dwarf galaxies has not been thoroughly explored, though some trends have been noted, such as AGN mildly suppressing central dark matter densities (see Koudmani et al., 2022; Arjona-Gálvez et al., 2024), hinting at a possible role of AGN in driving cusp-to-core transformations in dwarfs. However, the central density suppressions by the AGN in these simulations were mostly minor and did not change the qualitative nature of the central profiles. This is likely due to the effective equation of state (Springel & Hernquist, 2003) employed for modelling the interstellar medium (ISM) in these simulations, which leads to smoother star formation histories alongside less bursty AGN feedback at low redshifts, often operating as ‘maintenance-mode’ feedback (also see discussion in Koudmani et al., 2022). Recently, Arora et al. (2024) examined the role of BH feedback in core formation within the NIHAO (Wang et al., 2015) simulations and found negligible effects in low-mass systems, largely due to inefficient BH growth in this mass regime, constrained by the Bondi accretion scheme. In contrast, they found significant impacts of BH feedback on core formation in more massive galaxies, aligning with earlier studies that show efficient AGN activity can create cores in these galaxies (e.g. Martizzi et al., 2013; Peirani et al., 2017; Macciò et al., 2020).
@@
It is therefore timely to investigate whether efficient AGN activity can induce cusp-to-core transformations in dwarf galaxies and potentially account for the observed diversity in dwarf rotation curves. Clearly, within this context it is crucial to model the multi-phase nature of the ISM but also to consider the role of non-thermal components such as magnetic fields or cosmic rays (CRs) (e.g. Uhlig et al., 2012; Booth et al., 2013; Hanasz et al., 2013; Salem & Bryan, 2014; Pakmor et al., 2016; Farber et al., 2018; Holguin et al., 2019; Dashyan & Dubois, 2020; Martin-Alvarez et al., 2020; Steinwandel et al., 2020). In particular, it has been shown that CRs may drive core formation (e.g. Hopkins et al., 2020b; Martin-Alvarez et al., 2023) though their impact in the dwarf regime remains controversial as gas-rich mergers may re-establish cores in these systems (see discussion in Martin-Alvarez et al., 2023). What is more, it has been found that CRs may significantly boost the efficiency of AGN feedback (e.g. Su et al., 2021, 2024; Byrne et al., 2024; Wellons et al., 2023) by suppressing cold gas accretion more effectively than thermal heating or momentum injection. Therefore, it is important to also assess the role of CRs when investigating cusp-core transformations with AGN.
@@ 2.1 General set-up
Our simulation suite is based on a set of $3$ initial conditions (ICs) of dwarf haloes selected from the fire ICs set, spanning a range of halo masses in the ‘classical dwarf’ regime. We select two low-mass dwarfs: m10q and m10y, with $z=0$ halo masses of $M_{\mathrm{vir}}=8\times 10^{9}$ $\mathrm{M_{\odot}}$ and $M_{\mathrm{vir}}=1.4\times 10^{10}$ $\mathrm{M_{\odot}}$, respectively. Our third halo, m10z, is an intermediate-mass dwarf with $z=0$ halo mass $M_{\mathrm{vir}}=3.5\times 10^{10}$ $\mathrm{M_{\odot}}$. All of these ICs were generated using the music IC generator (Hahn & Abel, 2011) and have been thoroughly explored with the fire-2 model (Hopkins et al., 2018). In the standard fire-2 runs, m10q has a quiescent growth history and is relatively isolated, m10y forms early and has a large core, and m10z is an ultra-diffuse galaxy. All three of our selected galaxies are central galaxies.
To simulate these systems, we use the gizmo hydrodynamics+gravity code, employing the mesh-free finite mass method (Lanson & Vila, 2008b, a; Gaburov & Nitadori, 2011; Hopkins, 2015, 2017) that tracks constant-mass gas resolution elements in a Lagrangian fashion. Our simulations include both baryons and dark matter, with mass resolutions of $M_{\mathrm{gas}}=262$ $\mathrm{M_{\odot}}$ and $M_{\mathrm{DM}}=1303$ $\mathrm{M_{\odot}}$, respectively. The gravitational softening for the dark matter, stars and BH particles is set to $\epsilon_{\mathrm{soft}}=0.0625\ \mathrm{ckpc}\,h^{-1}$. Note that we keep the softening for these particles comoving throughout the simulation. The softening for the gas resolution elements is set adaptively. This allows us to easily resolve core formation as demonstrated in previous work (e.g. Chan et al., 2015; Hopkins et al., 2018).
@@ 2.2 Galaxy formation model
For the galaxy formation physics, we use the newly updated fire-3 model (Hopkins et al., 2023a) which tracks the state of the multi-phase ISM and multiple forms of stellar feedback, including feedback from SNe [^1] We couple the SN momentum following Hopkins et al. (2023a) and do not yet include the updates to the momentum coupling from Hopkins (2024). Though we note that these changes mainly affect high-mass galaxies at low resolution, whilst the impact on high-resolution dwarfs is very limited. of Type Ia and II, stellar winds from OB and AGB stars as well as multi-wavelength photo-heating and radiation pressure. Star formation occurs in gas that is molecular, dense, self-gravitating, self-shielding, and Jeans-unstable. All simulations include magnetic fields, using the magneto-hydrodynamics methods from Hopkins & Raives (2016) and Hopkins (2016). The fire-3 model also has new optional physics modules compared to the previous fire (Hopkins et al., 2014) and fire-2 (Hopkins et al., 2018) models. Two of the most important optional features are a novel sub-grid BH model, including seeding, accretion and feedback, as well as the inclusion of CR physics. To separate the impacts of CRs and BHs, we perform four simulations for each set of ICs, including runs without CRs or BHs, with both CRs and BHs and with only CRs or BHs, respectively. For the runs without BHs but with CRs, no BHs are seeded and the CRs only stem from SNe. For the runs without CRs and with BHs, we incorporate all AGN feedback channels except for CR injection and SNe do not inject CRs either. See Table 1 for an overview of the simulation suite. In this table, we also indicate the central dark matter densities of each simulation run measured at $r=150$ pc, a characteristic radius for distinguishing cusps from cores in observations (see Read et al., 2019). For reference, we further provide the ratio between the dark matter densities from the full MHD runs and the dark matter only runs and a qualitative assessment of the central dark matter density shape as weak/strong cusp or core. The latter categorisation is decided by comparing the central densities to the expectations from the fiducial NFW profile and a cored profile following the coreNFW profile from Read et al. (2016). Weak cusps fall below the mean relation for an NFW profile but within the $2\sigma$ scatter. Similarly, weak cores lie above the coreNFW expectations but within the $2\sigma$ scatter (see Figure 9).
@@
Table 1: Overview of dwarf zoom-in simulation suite. We list the simulation names (first column), ICs (second column), and whether BHs (third column) and CRs (fourth column) are included. We also provide the $z=0$ BH mass (where applicable, see fifth column), central dark matter densities at $r=150$ pc (sixth column), the ratios between the central dark matter density for the hydro runs and dark matter only simulations (seventh column), and the shape of the central dark matter density profile (eighth column) as shown in Figure 9. All masses and density estimators are rounded to two significant figures.
@@
Simulation name ICs BHs? CRs? $\bm{M_{\mathrm{BH}}\ [10^{4}\,\mathrm{M_{\odot}}]}$ $\bm{\rho_{\mathrm{DM,inner}}\ [10^{8}\ \mathrm{M_{\odot}}\,\mathrm{kpc^{-3}}]}$ $\bm{\rho_{\mathrm{DM,inner}}/\rho_{\mathrm{DMO,inner}}}$ Central dark matter profile
m10q m10q No No / 0.73 0.47 weak core
m10qAGN m10q Yes No 2.7 0.42 0.27 strong core
m10qCR m10q No Yes / 1.4 0.92 weak cusp
m10qCR+AGN m10q Yes Yes 0.5 1.5 0.99 weak cusp
m10y m10y No No / 0.20 0.062 strong core
m10yAGN m10y Yes No 9.3 0.19 0.058 strong core
m10yCR m10y No Yes / 0.83 0.26 weak core
m10yCR+AGN m10y Yes Yes 2.2 0.74 0.23 weak core
m10z m10z No No / 0.22 0.10 strong core
m10zAGN m10z Yes No 7.0 0.069 0.032 strong core
m10zCR m10z No Yes / 0.11 0.053 strong core
m10zCR+AGN m10z Yes Yes 1.7 0.66 0.31 weak core
@@
The modelling of BHs and CRs remains highly uncertain. Several studies have investigated the impact of different modelling assumptions for CRs (e.g. Hopkins et al., 2020b, 2023b; Butsky et al., 2023) and BHs (Anglés-Alcázar et al., 2017b, 2021; Su et al., 2021, 2024; Byrne et al., 2023, 2024; Wellons et al., 2023) within the fire model. For our study, we base our BH model on the extensive study by Wellons et al. (2023) who identified a modelling space that allows for efficient BH feedback whilst still matching observational constraints, in particular with regards to star formation distributions. We summarise these BH modelling choices in Section 2.2.1 and point the interested reader to Wellons et al. (2023) for more details and an in-depth parameter study of alternatives to our choices. For the CR modelling, we employ the numerically-efficient subgrid CR model introduced in Hopkins et al. (2023b) allowing us to explore several different IC set-ups to $z=0$ whilst minimising the computational overhead costs that would arise from explicit CR transport. We summarise the key characteristics of the CR subgrid model in Section 2.2.2 and refer the interested reader to Hopkins et al. (2023b) for a more detailed overview of this approach.
@@ 2.2.1 BH modelling
Once a BH particle is present, it grows via accretion by draining gas from the nearest $N\sim 256$ gas cell neighbours. Wellons et al. (2023) parameterized the accretion rate as $\dot{M}_{\mathrm{acc}}\equiv\eta_{\mathrm{acc}}M_{\mathrm{gas}}\Omega$, where $\eta_{\mathrm{acc}}$ is the accretion efficiency, $M_{\mathrm{gas}}$ is the gas mass within the BH kernel, and $\Omega=\sqrt{GM_{\mathrm{tot}}/R}$ is the orbital frequency determined at the kernel size $R$, based on the total mass enclosed within $R$, $M_{\mathrm{tot}}$. fire-3 modulates the BH accretion rate via a gas reservoir that represents an accretion disc from which the BH particle grows at a mass growth rate $\dot{M}_{\mathrm{BH}}=M_{\mathrm{\alpha disc}}/t_{\mathrm{dep}}$, where $M_{\mathrm{\alpha disc}}$ is the mass of the accretion disc and $t_{\mathrm{dep}}=42~{}\mathrm{Myr}\ (1+M_{\mathrm{BH}}/M_{\mathrm{\alpha disc}})^{0.4}$ is the depletion time.
@@
We set $\epsilon_{\mathrm{jet}}=0.35$ which leads to an effective base opening angle of $\sim 7\degree$. The angle weighting is applied with respect to the total gas angular momentum accreted by the BH, $\bm{J}_{\mathrm{gas}}$, which is measured in the centre-of-mass frame of the BH – accretion disc sink particle based on the kernel-weighted accreted mass and relative position and velocity of the respective gas cell neighbour. This allows for AGN energy injection without significantly disrupting the small-scale gas supply in the central region which favours repeated bursts and/or continuous AGN activity allowing us to probe the ‘physically interesting’ regime for AGN influencing cusp to core transformations. As discussed in Su et al. (2021), the large-scale collimation properties of the jet are not hugely sensitive to the small-scale opening angle since the feedback energy chooses the path of least resistance effectively recollimating even for large opening angles, also see Koudmani et al. (2019). Long-range radiation transport and CRs are treated via the LEBRON approximation and injected isotropically (see Hopkins et al., 2020a, 2023b).